OpenStudy (anonymous):

-x - 4y=-37 -2x -5y=-53

OpenStudy (anonymous):

ok, so now: multiply the first equation by -2 to get 2x+8y=74. Add this equation to the second equation to get: 3y=21. Solve for y to get y=7. Substitute 7 into the first equation to get: -x-4(7)=-37 -x-28=-37 -x=9. To get positive x, divide by -1. x=-9.

OpenStudy (anonymous):

Confused with anything? let me know!

OpenStudy (anonymous):

thankssss sweetie soo much =) muchos GRAcias!!

OpenStudy (anonymous):

no problem at all. Denada!

OpenStudy (anonymous):

wait it was wrong =(

OpenStudy (anonymous):

yea right!

OpenStudy (anonymous):

oh i think i found it.

OpenStudy (anonymous):

stupid negatives :( -x-28=-37. correction: -x=-9. so x=9. SOOOO SORRYYY!!!

OpenStudy (anonymous):

well can u help me again ???? and get it right this time? please

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

3x=9 - 6y 3x+8y=9

OpenStudy (anonymous):

replied to it on the new post.